OpenStudy (anonymous):

what are the zeros of the function f(x) = x^4 - 4x^3 + 3x^2 ? what are their multiplicities?

OpenStudy (jim766):

factor out the gcf x^2

OpenStudy (jhannybean):

as what @Jim766 said, the problem would be easier to solve. \[\large x^2(x^2-4x+3)=0\] Now concentrate on factoring just the \[\large x^2-4x+3=0\] first. what two numbers add to give -4 but multiply to give +3?

OpenStudy (anonymous):

-1 and -3

OpenStudy (jhannybean):

good :) so we can use -1 and -3 to factor out x^2 -4x+3 \[\large (x^2-4x+3) \implies (x-3)(x-1)=0\] so now you have \[\large x^2=0\]\[\large x-3=0\]\[\large x-1 = 0\] can you solve for all the x's? :)

OpenStudy (anonymous):

x = 1 x = 3 and x = 0

OpenStudy (jhannybean):

good job!! There you go :D

OpenStudy (jhannybean):

You found the zeroes of the function !

OpenStudy (anonymous):

are those the multiplicities too?

OpenStudy (jhannybean):

mmhmm :)

OpenStudy (anonymous):

great. thank you :)

OpenStudy (jhannybean):

no problem :)

OpenStudy (precal):

multiplicity refers to their powers

OpenStudy (precal):